OpenStudy (anonymous):

How would you solve? log x=log(x^2-12) ?

OpenStudy (anonymous):

\[\log x=\log(x^2-12)\rightarrow x=x^2-12\rightarrow x^2-x-12=0=(x+4)(x-3)\]So x=3 (can't be -4 due to domain restrictions on the log function)

OpenStudy (anonymous):

Actually, neither solution is any good. No solution.

OpenStudy (anonymous):

Luis, you did the same calculation I did, but neither of those solutions meet the specifications of the problem. If x=3, the RHS has a negative number in the argument of the log. If x=-4, the same problem on the LHS. No solution.

OpenStudy (anonymous):

you didn't factor x^2-x-12 properly. the answer is x=4

OpenStudy (anonymous):

luis did it right

OpenStudy (anonymous):

we know -4 + 3 is -1 and -4 * 3 is -12 so X^2-X-12=X^2 -4x+3x- 4 *3 = X(X-4)+3(X-4) =(X-4)(X+3)

OpenStudy (anonymous):

Yep. Epic fail.................